OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (jdoe0001):

just like the ones before, test each choice, see which one gives true for both inequalities

OpenStudy (anonymous):

What do you mean? @jdoe0001

OpenStudy (jdoe0001):

we have done these ones before about 2 or 3 days ago IIRC

OpenStudy (jdoe0001):

say for example testing the 1st point (1, 10) x y (10) <= (1)-5 => 10 <= -4 so is 10 really less than -4? well that's not true, thus is not a solution and then test the other points

OpenStudy (anonymous):

I don't understand what you just did. @jdoe0001

OpenStudy (jdoe0001):

hmm ----> y less than or greater to x - 5 is this supposed to be -> y less than or EQUAL to x - 5?

OpenStudy (anonymous):

Equal to.

OpenStudy (jdoe0001):

so in the system of equations \(\bf y \le x-5\\ y \le -x-4\) you use the 2nd point and test, and then the 3rd and test and then the 4th one the solution will be where the inequalities become true as you see above, I replace my (x, y) = (1, 10), so x = 1, y =10 in the 1st equation I ended up with 10 <= -4 but of course 10 is not less than -4, thus is not true thus is not a solution

OpenStudy (anonymous):

So you replaced (x, y) = (1, 10) and you got 10 <= -4?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

of course 10 is not less than -4, thus is not true thus is not a solution

OpenStudy (jdoe0001):

so the solution point, will give TRUE on both equations when replacing the (x,y) for the given point something TRUE will be say 3 < 6, or 2 < 5, or -1 > -5 or 100 > 99 something FALSE will be say 1 > 2, or 3 < 1 and such

OpenStudy (jdoe0001):

they're false because 1 is not greater than 2, and 3 is not less than 1

OpenStudy (jdoe0001):

3< 6 is true because, indeed, 6 is really greater than 3, by 3 units and 3 is indeed smaller than 6

OpenStudy (anonymous):

Yeah I know that but how exactly do you solve this? You replace (x, y) = (1, 10) but then what?

OpenStudy (jdoe0001):

(1, 10) x y (10) <= (1)-5 => 10 <= -4 then expand and add each side

OpenStudy (anonymous):

How do you expand and add each side?

OpenStudy (jdoe0001):

same way you'd simplify any terms

OpenStudy (jdoe0001):

all I did was replace the x and y for the values given in the point and use those values in the x and y in the equation

OpenStudy (anonymous):

Can you show me that in equation form instead of words?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

(1, 10) x y \(\bf y \le x-5\) (10) <= (1)-5 => 10 <= -4

OpenStudy (anonymous):

Could you explain how you got the second part of that last line?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

=> 10 <= -4

OpenStudy (jdoe0001):

\(\bf (\color{blue}{10}) <= (\color{green}{1})-5 => \color{blue}{10} <= \color{green}{1}-4\)

OpenStudy (jdoe0001):

meh

OpenStudy (jdoe0001):

\(\bf (\color{blue}{10}) <= (\color{green}{1})-5 => \color{blue}{10} <= \color{green}{1}-5\)

OpenStudy (jdoe0001):

1-5 = -4

OpenStudy (anonymous):

Could you explain it?

OpenStudy (jdoe0001):

hmmm what's confusing you?

OpenStudy (anonymous):

(10)<=(1)−5=>10<=1−5 What are you doing there?

OpenStudy (jdoe0001):

\(\bf (\color{blue}{10}) \le (\color{green}{1})-5 \implies \color{blue}{10} \le \color{green}{1}-5\)

OpenStudy (jdoe0001):

just adding, nothing else

OpenStudy (anonymous):

Adding the whole thing?

OpenStudy (jdoe0001):

nope adding each side by itself

OpenStudy (anonymous):

And how do you do that? What exactly are you adding?

OpenStudy (jdoe0001):

if the expression were for example lemme give you a different inequality and point y + 2y <= 3x -2x+5 point will be ( 5, 7 ) x y (7) + 2(7) < = 3(5)-2(5)+5 # all y = 7, x = 5 thus 7 + 14 <= 15-10+5 21 <= 10 now is 21 really less than 10? well then is a false inequality that means the point (5, 7) is not a solution for that inequality

OpenStudy (anonymous):

# all y = 7, x = 5 thus What?

OpenStudy (jdoe0001):

y = 7, x = 5

OpenStudy (jdoe0001):

is all

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (jdoe0001):

yw